An order was written for nystatin (Mycostatin) 750,000 units to be given three times a day...

  1. An order was written for nystatin (Mycostatin) 750,000 units to be given three times a day with the instructions to swish and swallow. On hand is nystatin (Mycostatin) 100,000 units/mL. How many teaspoons do you prepare per dose?
Answer #1

Ordered Nystatin= 750,000 units to be given three times a day.

On hand available= 100,000 units/ ml

For the order required ml= 750,000/100,000 ml

=7.5 ml three times a day

5 ml= teaspoon

7.5 ml = 7.5/5= 1.5 teaspoon should be used per dose to fulfill the order.